# Task Facet Learning: A Structured Approach to Prompt Optimization

### v1

@@ Abstract
Given a task in the form of a basic description and its training examples, prompt optimization is the problem of synthesizing the given information into a text prompt for a large language model (LLM). Humans solve this problem by also considering the different facets that define a task (e.g., counter-examples, explanations, analogies) and including them in the prompt. However, it is unclear whether existing algorithmic approaches, based on iteratively editing a given prompt or automatically selecting a few in-context examples, can cover the multiple facets required to solve a complex task. In this work, we view prompt optimization as that of learning multiple facets of a task from a set of training examples. We identify and exploit structure in the prompt optimization problem — first, we find that prompts can be broken down into loosely coupled semantic sections that have a relatively independent effect on the prompt’s performance; second, we cluster the input space and use clustered batches so that the optimization procedure can learn the different facets of a task across batches. The resulting algorithm, UniPrompt, consists of a generative model to generate initial candidates for each prompt section; and a feedback mechanism that aggregates suggested edits from multiple mini-batches into a conceptual description for the section. Empirical evaluation on multiple datasets and a real-world task shows that prompts generated using UniPrompt obtain higher accuracy than human-tuned prompts and those from state-of-the-art methods. In particular, our algorithm can generate long, complex prompts that existing methods are unable to generate. Code for UniPrompt will be available at https://aka.ms/uniprompt.
@@ 1 Introduction
Given a task, choosing an input prompt is a key part of optimizing Large Language Model’s (LLM) performance [14, 31]. Minor changes in prompt can lead to performance gains or losses, necessitating prompt engineering [16] for effective model utilization. Typically, manually-developed prompts combine task description with a few in-context examples, along with modifiers like chain-of-thought [14]. For greater accuracy, human prompt engineers spend considerable time to identify errors with a current prompt, consider the different facets of a task (e.g., counter-examples, explanations, analogies) that may fix those errors, and include them in the prompt. For instance, for a hate speech classification task, in addition to the definition, it may be helpful to specify the facets that lead to hate speech: the context of conversation, identifying intent, and differentiating hate speech from opinions or closely-related concepts such as vulgarity and profanity.
@@
To avoid the cumbersome process of manually creating prompts, recent work aims to automate the process of generating natural language prompts that are also interpretable. Since language tokens are discrete, this leads to a challenging discrete optimization problem with a combinatorial space of possible outputs. Techniques for prompt optimization can be divided in two categories: non-directional, e.g., random search [36, 35] and genetic algorithms [31, 7], where the sampling of new input is “random” and does not explicitly aim to reduce error on a train set; and directional, where the sampling of new input depends on some error measure on a representative train sample. Recently, more complex methods are proposed in the second category including reinforcement learning [33, 6] and updating prompts using auxiliary LLMs to reduce incorrect predictions over a train sample [12, 22]. While all these techniques focus on editing, adding, or deleting parts of a given prompt, they are developed with the goal of obtaining a concise description of the task. None of them focus on ensuring that multiple facets of a task are added to the prompt.
In this paper, we study how to optimize prompts to cover diverse, multiple facets of a task and improve overall accuracy. To develop a suitable algorithm, we analyze properties of the prompt optimization problem. First, we check how sensitive LLMs are to small changes in the input prompt, since directional optimization algorithms may not be as useful if LLMs are highly sensitive. Using a notion of probabilistic Lipschitz continuity, we find that LLMs become less sensitive with size. Across prompt edits with the same semantic meaning, Llama-13B accuracy fluctuates up to 63% of the minimum accuracy whereas the relative accuracy change in GPT-3.5 and GPT-4 models is 6% and 4% respectively. As a result, probabilistic Lipschitz constant for GPT-3.5 and GPT-4 is significantly smaller than that for Llama-13B, indicating that directional optimization methods may be best suited for larger models. Second, we study how LLM accuracy changes as more information is added to the prompt. Ideally, we would want a diminishing marginal returns (sub-modularity) property so that greedy addition of different facets can lead to a reasonable approximation of the true optimum. We find that diminishing returns property largely holds when adding different facets as new sections to the prompt, but it does not hold when directly adding in-context examples.
Our analysis provides two insights: 1) directional optimization may be feasible for models like GPT-3.5 having a low probabilistic Lipschitz constant; 2) including facets as additional text sections may lead to a well-behaved optimization problem. Based on these, we propose an algorithm, UniPrompt, that can generate the full prompt from scratch, given a one line description for a task. Specifically, we identify a prompt structure that provides different task facets as sections in the prompt and initializes them with text generated from a language model. To develop a high-accuracy prompt, we use an auxiliary LLM to obtain feedback about example predictions with the current prompt and decide to make edits at a section level (add, edit or delete) in each iteration. Our key insight is that feedback on a single example or a randomly selected batch of examples does not yield generalizable facet descriptions. Instead, we propose generating edits at a minibatch level and aggregating the edits at the batch level to yield a single edit that conveys a broad concept relevant to the task. To further increase the chances of facet learning, we cluster the input examples beforehand and use clustered mini-batches (see Figure 1).
@@
We evaluate UniPrompt on zero-shot (no in-context examples) and few-shot settings on four different tasks. In the zero-shot setting, UniPrompt consistently achieves higher accuracy than existing prompt optimization methods. On Ethos, a hate speech dataset, UniPrompt obtains 94% accuracy whereas the next best method obtains 82%. In the few-shot setting, we compare UniPrompt to MedPrompt [21], a state-of-the-art prompt composition method. We find that UniPrompt, requiring only one LLM call at inference time, obtains the same accuracy as MedPrompt that requires five calls. If we allow multiple calls to UniPrompt, we obtain over 4% accuracy gains. We also evaluate UniPrompt on a real-world semantic matching task in a web search engine. Compared to the best manual prompt, the prompt generated from UniPrompt leads to over 5% increase in accuracy on the negative class and nearly 2% accuracy increase overall.
@@
Figure 1: Existing prompt optimization methods (left) versus UniPrompt (right) on the SciQ [29] dataset: [Left] State-of-the-art prompt optimization methods such as ProTeGi  [22] sample from the questions wrongly answered by the current prompt, and use an expert LLM (such as GPT-4) to obtain feedback on the mistakes made; This approach tends to overfit to specific examples, as can been seen from the proposed prompts $p^{\prime}$; leading to limited generalization. [Right] In contrast, UniPrompt identifies key task facets that generalize across examples by: (1) clustering examples into groups with similar task facets, and (2) employing a two-tier feedback-based update strategy. The resulting prompt updates extract generalizable concepts from the specific examples seen.
2 Characterizing the Prompt Optimization Problem
Assume access to black-box solver LLM $f:\mathcal{X}\rightarrow\mathbb{R}$ that takes as input a prompt $\mathbf{x}\in\mathcal{X}$ and outputs the average accuracy on a validation set $D_{v}$. Since the set of prompts is combinatorially large, we assume that all prompts can be embedded in a fixed $p$-dimensional space such that distance between two prompts in the vector space correspond to their semantic similarity. Thus, each prompt $\mathbf{x}$ can be expressed as a vector in $\mathbb{R}^{p}$. The prompt optimization problem can be written as $\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x};D_{v})$.
@@ 2 Characterizing the Prompt Optimization Problem
2.1 Is directional optimization feasible?
Previous work has shown that LLMs can be brittle to their input: changing the prompt slightly can create a significant difference in performance [37]. To understand whether optimization algorithms may work, it is important to understand how the surface of $f$ looks. If $f$ is brittle everywhere to small changes, then random search over prompts (such as in [36]) may be a competitive method; it is unclear whether optimization can help. On the other hand, if the change in $f$ given small changes to the input prompt is bounded, then optimization algorithms can hope to perform significantly better than random search.
More formally, we ask whether the prompt optimization problem is well-conditioned. Typically, conditioning can be determined by the Hessian. However, since $f$ is assumed to be black-box, we approximate it by measuring sensitivity, or more specifically, Lipschitz continuity near the optimal solution. Based on prior work on defining continuity of neural networks [18], we use a probabilistic notion.
Definition 1
Probabilistic Lipschitz Continuity [18]. Given a probability distribution over inputs $\mathcal{X}$, $r\geq 0$, a function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $(L,\epsilon)$-probabilistically Lipschitz with constant $L\geq 0$, if
$\Pr_{\mathbf{x},\mathbf{x}^{\prime}\sim\mathcal{X}}[\operatorname{d}(f(\mathbf{x}),f(\mathbf{x}^{\prime}))\leq L.\operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\ |\ \operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\leq r]\geq 1-\epsilon$ (1)
where $d$ is a distance measure such as $\ell_{1}$ or $\ell_{2}$ norm.
Note the focus on small changes in input through the parameter $r$. Intuitively, the Lipschitz property bounds the maximum change in $f$ given a small change in input prompt. Typically, the lower bound of error for any sequential optimization algorithm over $f$ is directly proportional to the Lipschitz constant $L$ [17]. Therefore, for faster convergence, it is desirable to have a low $L$, especially near the optimal solution. Empirically, we estimate $L$ by sampling task-relevant prompts so that they are close to the optimal solution. Then we make small changes to the prompt such that the semantic meaning stays the same and measure the change in $f$. We show the change in $f$ per change in input for GPT-4, GPT-3.5, and Llama2-13B models in Figure 2(a) for the Ethos dataset commonly used in prompt optimization [22, 19, 32, 23, 10]. Assuming $\epsilon=0.05$, probabilistic Lipschitz constant $L$ for GPT-3.5 and GPT-4 is $<1$, whereas it is over three times more for Llama2-13B. Thus, as the size of the model increases, the probabilistic Lipschitz constant decreases, indicating that larger models like GPT-3.5 and GPT-4 are more robust to small changes and thus amenable to optimization algorithms. For Llama, the absolute change in accuracy ranges between 0.33 to 0.54, indicating a 63% change in accuracy even across prompts with the same semantic meaning. In contrast, the relative change in accuracy for GPT-3.5 and GPT-4 is 6% and 4% respectively. (See Appendix A.1 for experimental details)
Figure 2: On the Ethos dataset, (a) estimating (probabilistic) Lipschitz constant of models (Definition 1); (b) diminishing marginal utility (submodularity) of GPT-3.5 on two types of prompting strategies.
2.2 Can higher-order abstractions of task examples help prompt optimization?
Consider in-context learning [3], where the prompt $\mathbf{x}$ is composed of a simple task description and a set of labeled examples from the task. The optimal prompt should consider all the aspects of the task which can only be discovered by exhaustively looking at the examples. But, in practice, we are constrained by (a) availability of a limited number of training examples, (b) prompt size. It is beneficial for in-context learning to consider the marginal utility of examples [34, 8]. For instance, it would help to add examples where the model fails rather than ones where the model already succeeds (see Table 4 in Appendix). This suggests that one could use a greedy algorithm for iteratively optimizing the prompt by finding failing examples and adding to the prompt. But would such a procedure converge to a good solution?
We can view $f(\mathbf{x})$ as a set function over the examples (that compose $\mathbf{x}$). A natural property to examine is submodularity (where greedy optimization provably converges to a good solution under some conditions [15]), i.e., the property of diminishing marginal utility of examples. The marginal utility of adding an example $e$ to the few-shot prompt corresponding to a set of examples $A$ is $\Delta_{A}=f(A\cup\{e\})-f(A)$. Let $B\subset A$ denote a subset of $A$. We define $f_{\text{fewshot}}$ to be the validation accuracy of GPT-3.5 using few-shot prompts on the Ethos dataset (see Section 4). In Figure 2 (b), we show the distribution of diminishing marginal utilities for $f_{\text{fewshot}}$ (light blue), i.e., we sample multiple $A,B$ sets of examples from the dataset and compute $\Delta_{B}-\Delta_{A}$, which is expected to be non-negative for submodular functions. As we see, the marginal utility differences are nearly uniformly random in $[-0.01,0.01]$, indicating that there is no submodularity structure to $f_{\text{fewshot}}$.
A recurring theme [1, 2] in the prompts that humans engineer is to organize task information in “sections” (introduction, context, etc.), treating prompt like a structured document. When humans engineer a prompt, they do it systematically by editing specific sections of the prompt. We hypothesize that there are a small number of facets that collectively define a task, which can be discovered by inspecting a representative set of examples from the data distribution. Consider the task of tagging hate speech. Identifying intent to harm, recognizing general aggression versus targeted hate, and differentiating hate from other objectionable content are some facets of the task. We ask whether there is submodularity structure in adding different facets (text sections) to a prompt, compared to adding in-context examples. The answer to this question depends on multiple factors: (1) what are the contents of the sections and how are they generated?; (2) how does $f(\mathbf{x})$ behave w.r.t. the sections that make up the prompt $\mathbf{x}$?
Figure 2 (b) shows the distribution of marginal utility differences for $f_{\text{facets}}$ (dark blue), i.e., we sample multiple $A,B$ sets of facets (i.e., short sections) that are generated using a Llama2-13B model, and compute $\Delta_{B}-\Delta_{A}$ (defined as above, but the added element $e$ is now a section) using GPT-3.5 as the solver LLM (See Appendix A.2 for experimental details). We see that the diminishing marginal returns property largely holds for $f_{\text{facets}}$, unlike in the case of $f_{\text{fewshot}}$. This suggests that greedy strategies for prompt optimization problem over sections could work — but this crucially relies on how we generate the sections themselves.
The prompt optimization problem can now be posed as one of inferring the facets of a task from a set of training examples. In the following section, we propose a method that iteratively learns the facets of the task as sections of the prompt.
3 UniPrompt: Generating a Prompt that Captures Multiple Facets of a Task
Manual prompt engineering consists broadly of two steps: 1) generating an initial prompt; and 2) identifying errors and editing the prompt. The initial prompt for a task is constructed based on domain knowledge. Typically, the initial prompt contains implicit or explicit sections (e.g., introduction, task description, intended application, edge-cases, challenges, etc.) wherein each section corresponds to a relevant task facet. Given the initial prompt and a solver LLM, a human prompt engineer would look at the incorrect predictions, form hypotheses on which section of the prompt needs to be edited or whether a new section needs to be added, and then propose edits that may fix the errors. Then, the prompt engineer may pick the edit with the highest validation accuracy. These steps are repeated in a loop, until the validation accuracy saturates. A good engineer may also backtrack from an optimization trace and start from some previous prompt, in case the current optimization trajectory yields unsatisfactory results.
We present UniPrompt, a prompt optimization algorithm that aims to mimic the above process. Specifically, the algorithm extracts key concepts or facets relevant to a task and updates prompt sections using them, with the overall goal of increasing validation accuracy. Our analysis from Section 2 indicates that such a directional optimization procedure based on iteratively adding task facets can converge to a good solution, especially for solver LLMs such as GPT-3.5 and GPT-4.
[…]
1) Task Facets Initialization using background knowledge. In the first stage, the algorithm leverages a language model to generate a prompt with multiple facets, i.e., sections focusing on specific task aspects based on the task description. It greedily selects content for each facet from a pool of candidates generated using the language model.
2) Task Facets Refinement using examples. In the second stage, the algorithm refines the prompt using a set of training examples and an auxiliary LLM. Given a candidate prompt, it collects examples with prediction errors by the solver LLM, extracts key concepts or facets required to fix those errors, and makes appropriate edits. An edit may create a new section, modify an existing section, or delete an existing section. Additionally, we maintain a beam of best performing prompts to backtrack and restart from a previous prompt if the current optimization path is unsatisfactory.
@@ 3 UniPrompt: Generating a Prompt that Captures Multiple Facets of a Task
Similar to UniPrompt, previous methods [12, 22] propose using an auxiliary LLM to provide feedback on incorrect predictions and propose edits to a given prompt. In our initial experiments, however, we find that collecting feedback over individual examples or randomly sampled batches leads to memorization of individual examples (see Figure 1) rather than recognition of facets that are important to the task. To avoid overfitting to the train examples and to extract generalizable concepts, we make two contributions. First, we follow a two-tier setup of synthesizing feedback for a batch of training examples. We break up a batch into mini-batches, collect feedback on each of the mini-batches and then use a separate prompt to aggregate the different feedback texts into a generalizable concept. Second, to increase chances that a mini-batch corresponds to a coherent facet, we cluster the training data beforehand and ensure that each mini-batch consists of examples from the same cluster. We provide details of the algorithm below (see Algorithm 1).
Notation: We denote the training set with $D_{t}$ where each example is a question-answer pair of the form $\{q_{i},a_{i}\}$. The validation set is denoted as $D_{v}$. We have $N$ training examples and $K$ validation examples. Input to the algorithm is the solver LLM $f_{LLM}$, train set $D_{t}$, validation set $D_{v}$, initial prompt for the task $p_{0}$, one-line task description $T$. In addition, we assume access to an “expert” auxiliary LLM such as GPT-4.

### v2

@@ Abstract
Given a task in the form of a basic description and its training examples, prompt optimization is the problem of synthesizing the given information into a text prompt for a large language model. Humans solve this problem by also considering the different facets that define a task (e.g., counter-examples, explanations, analogies) and including them in the prompt. However, it is unclear whether existing algorithmic approaches, based on iteratively editing a given prompt or automatically selecting a few in-context examples, can cover the multiple facets required to solve a complex task. In this work, we view prompt optimization as that of learning multiple facets of a task from a set of training examples. We exploit structure in the prompt optimization problem and break down a prompt into loosely coupled semantic sections. The proposed algorithm, UniPrompt, (1) clusters the input space and uses clustered batches so that each batch likely corresponds to a different facet of the task, and (2) utilizes a feedback mechanism to propose adding, editing or deleting a section, which in turn is aggregated over a batch to capture generalizable facets. Empirical evaluation on multiple datasets and a real-world task shows that prompts generated using UniPrompt obtain higher accuracy than human-tuned prompts and those from state-of-the-art methods. In particular, our algorithm can generate long, complex prompts that existing methods are unable to generate. Code for UniPrompt is available at https://aka.ms/uniprompt.
Task Facet Learning: A Structured Approach To Prompt Optimization
Gurusha Juneja1,*, Gautam Jajoo2,*, Nagarajan Natarajan3, Hua Li4, Jian Jiao4, Amit Sharma3,† 1UC Santa Barbara, 2BITS Pilani, 3Microsoft Research, 4Microsoft Bing Ads *These authors contributed equally to this work. Corresponding author: amshar@microsoft.com
55footnotetext: To appear in ACL Findings 2025.
@@ 1 Introduction
Given a task, choosing an input prompt is a key part of optimizing Large Language Model’s (LLM) performance (Kojima et al., 2024; Yang et al., 2023). Minor changes in prompt can lead to performance gains or losses, necessitating prompt engineering (Liu et al., 2023). Typically, manually-developed prompts combine task description with a few in-context examples, along with modifiers like chain-of-thought (Kojima et al., 2024). For greater accuracy, human prompt engineers spend considerable time to identify errors with a current prompt, consider the different facets of a task (e.g., counter-examples, explanations, analogies) that may fix those errors, and include them in the prompt. For instance, for a hate speech classification task, in addition to the definition, it may be helpful to specify the facets that lead to hate speech: the context of conversation, identifying intent, and differentiating hate speech from opinions or closely-related concepts such as vulgarity and profanity.
@@
To avoid the above cumbersome manual process, recent work aims to automate the process of generating natural language prompts that are also interpretable. Since language tokens are discrete, this leads to a challenging discrete optimization problem with a combinatorial space of possible outputs. Techniques for prompt optimization can be divided in two categories: non-directional, e.g., random search (Zhou et al., 2022; Zhang et al., 2023) and genetic algorithms (Yang et al., 2023; Guo et al., 2023), where the sampling of new input is “random” and does not explicitly aim to reduce error on a train set; and directional, where the sampling of new input depends on some error measure on a representative train sample. Recently, more complex methods have been proposed in the second category including RL (Zhang et al., 2022; Deng et al., 2022), updating prompts using feedback from auxiliary LLMs (Hu et al., 2024; Pryzant et al., 2023), and optimizing the input to a small LM that generates the prompt (Lin et al., 2024b; Chen et al., 2024). While these techniques focus on editing parts of a given prompt, they are developed with the goal of obtaining a concise description of the task. None of these focus on ensuring multiple facets of a task are added to the prompt.
In this paper, we propose UniPrompt, a prompt optimization method to cover diverse, multiple facets of a task and improve overall accuracy. To simulate the manual prompt engineering process, we propose that prompts be constructed from individual sections, where each section may correspond to a different facet that humans may consider for the task. Prompt editing proceeds at a section-level: we can add, edit or delete a section from the prompt. Similar to Pryzant et al. (2023); Hu et al. (2024), prompt edits are based on an auxiliary LLM’s feedback about example predictions with the current prompt. We contribute two key insights in this feedback-based optimization process. First, we find that the feedback on a single example or a randomly selected batch of examples does not yield generalizable facet descriptions. Instead, we propose clustering the inputs and creating mini-batches such that each mini-batch is sourced from a single cluster. Second, even with clustered batches, the feedback tends to overfit to specific examples or their properties. To generate a prompt edit that conveys a generalizable concept relevant to the task, we propose generating edits at a mini-batch level and then aggregating them at the batch level to yield the final edit (Figure 1). While the two insights may appear simple, we show that they significantly improve extracting diverse task facets.
@@
We evaluate UniPrompt on several benchmarks where it consistently achieves higher accuracy than existing prompt optimization methods. On Ethos, a hate speech dataset, UniPrompt obtains 94% accuracy whereas the next best method obtains 82%. Even though UniPrompt focuses only on the instruction and does not include any in-context examples, we find that its instruction-only accuracy is often higher than methods such as DSPy (Khattab et al., 2024) that optimize both. In the few-shot setting, we also compare UniPrompt to MedPrompt (Nori et al., 2023), a state-of-the-art prompt composition method. We find that UniPrompt, requiring only one LLM call at inference time, obtains the same accuracy as MedPrompt that requires five calls. If we allow multiple calls to UniPrompt, we obtain over 4% accuracy gains. Finally, we also evaluate UniPrompt on a real-world semantic matching task in a web search engine. Compared to the best manual prompt, the prompt generated from UniPrompt leads to over 5% increase in accuracy on the rare class and nearly 2% accuracy increase overall.
@@
Figure 1: Existing prompt optimization methods (left) versus UniPrompt (right) on the Ethos dataset: [Left] State-of-the-art prompt optimization methods like ProTeGi  (Pryzant et al., 2023) sample from the questions wrongly answered by the current prompt, and use an expert LLM (e.g., GPT-4) to obtain feedback on the mistakes. This approach tends to give very general edits or overfits to specific examples. [Right] In contrast, UniPrompt identifies key task facets by: (1) clustering examples with similar task facets, and (2) employing a two-tier feedback-based update strategy. The resulting prompt updates extract generalizable concepts from the specific examples.
2 Related Work
Here, we highlight relevant work that are not addressed in the manuscript otherwise. Deng et al. (2022) present a discrete prompt optimization method, RLPrompt, using reinforcement learning, where a policy network learns to generate effective prompts through reward-based training, with an emphasis on enhancing training efficiency through effective reward stabilization techniques. A drawback of such automatic prompt optimization approaches (Pryzant et al., 2023; Zhou et al., 2022; Deng et al., 2022; Yang et al., 2023) is that the prompts generated tend to be short, often comprising only one or two sentences, which may not fully encapsulate the complexity of the task at hand.
Another recent line of work leverages human feedback. Automated Prompt Optimization with Human Feedback  (Lin et al., 2024a) optimizes prompts for black-box LLMs using human preference feedback. Besides the obvious overhead, it might also introduce potential biases.
Prior research (Wei et al., 2023, 2024) has highlighted the significance of specific sections within prompts. However, existing methods do not specifically target the optimization of individual sections and their respective contents within the prompts. Hsieh et al. (2023) investigate the use of greedy and genetic algorithms to edit lengthy prompts. Their method focuses on paraphrasing one line at a time starting from an existing prompt, compared to our goal of learning facets of a task from scratch. Another orthogonal line of work explores algorithmic selection of in-context examples (Min et al., 2022; Gupta et al., 2023; Wu et al., 2023; Srivastava et al., 2024; Sun et al., 2024).
3 UniPrompt: Capturing Task Facets
@@
1. Larger models are more amenable to prompt optimization. We observe that the change in the objective function (i.e., loss on a validation set for a given prompt) per change in input is relatively more stable for larger models like GPT-4 (Figure 2) than for GPT-3.5 (analysis in Appendix A.1).
2. Clustered-batching improves the quality of text gradients (i.e., feedback), as against the standard random batching adopted in state-of-the-art prompt optimization methods (Section 5.1).
3. Two-tier feedback helps learn generalizable facets. Collecting feedback from an expert LLM over mini-batches, and then summarizing the individual feedback texts via a second step (Section 5.1) helps learn generalizable task concepts in prompts.
@@
The proposed method UniPrompt, in Figure 1 [Right], makes two contributions. First, we follow a two-tier setup of synthesizing feedback for a batch of training examples. We break up a batch into mini-batches, collect feedback on each of the mini-batches and then use a separate prompt to aggregate the different feedback texts into a generalizable concept. Second, to increase chances that a mini-batch corresponds to a coherent facet, we periodically (re)cluster the training data and ensure that each mini-batch consists of examples from the same cluster.
Algorithm 1 receives as input a one-line task description and a train set $D_{t}$ of $N$langlequestion $q_{i}$, answer $a_{i}\rangle$ demonstrations. It extracts key concepts or facets relevant to the task and updates prompt sections using them, with the goal of increasing accuracy on the validation set $D_{v}$. We assume access to an “expert LLM” such as GPT-4.
3.1 Task facet learning using examples
Extracting task-relevant concepts from a set of examples to refine a prompt is a complex problem comprising multiple steps. Given a set of incorrect predictions, one needs to analyze what went wrong in each prediction, form hypotheses, aggregate the hypotheses to identify specific concepts that are relevant for the task. Then, for each concept, one needs to attribute which facet/section of the current prompt needs to be edited to incorporate the concept. These operations are highly model-specific and are difficult to execute reliably. Therefore, we exclusively rely on an expert LLM.
First, we prompt the expert LLM to diagnose mistakes (feedback) in each example given the answer and chain-of-thought reasoning produced by the solver LLM. Subsequently, we use this feedback to generate precise edits for the prompt that may fix the error. These individual edits are then aggregated over a mini-batch and fed back into the same LLM, which then identifies a few major edits to be applied to the current prompt. To aid in identifying major edits that correspond to generalizable facets, we propose to cluster the examples as a preprocessing step and create clustered batches, such that each cluster shares some common facet of the task.
3.1.1 Clustering for identifying facets
We explore two approaches for clustering: topic-based clustering, and feedback-based clustering.
Topic-Based Clustering. Given a set of examples, we identify $l$ topics spanning the entire train set. This type of clustering is motivated by the observation that solver LLM may make similar mistakes on examples from the same topic. Hence, for such examples, a common edit to the prompt could improve predictions for all the examples. To obtain the clusters, the expert LLM is prompted (for prompt see Appendix A.10) to provide a broad sub-topic $t_{i}$ for each question. Then the resultant list of sub-topics $\{t_{1},t_{2},\dots,t_{N}\}$ is again clustered into $k$ topics $\{t^{\prime}_{1},t^{\prime}_{2},\dots,t^{\prime}_{l}\}$ by prompting the expert LLM. Based on this clustering, each example ${q_{i},a_{i}}$ is assigned a cluster $t^{\prime}_{j}$ corresponding to $t_{i}$.
Feedback-Based Clustering. Examples that receive similar feedback based on a prompt’s predictions can help identify task facets. Consider a physics-based task where two examples from different topics obtain the same feedback from the expert LLM to edit the “Rules” section of the prompt to include the statement, “Draw all forces on each body before writing the equations”. We argue that such examples can be clustered. This type of clustering makes the broad edit identification step easier. To obtain the clusters, we first evaluate all training examples against the current best prompt and store the feedback $f_{i}$ from the expert LLM, corresponding to each incorrectly answered example ${q_{i},a_{i}}$ (all the correctly answered questions form one cluster). We then prompt the expert LLM to cluster these feedbacks $\{f_{1},f_{2},\dots,f_{N}\}$ into $l$ clusters (see Appendix A.11). For each cluster, we create a batch ${q_{i},a_{i}}$ corresponding to feedbacks in that cluster.
3.1.2 Obtaining generalizable prompt edits
Two-tier Feedback. To encourage generalizable feedback from the expert LLM, we obtain feedback at two levels: mini-batch and batch. Given a batch (created using clustering discussed above), we break it up into mini-batches.
For each mini-batch $m$, we construct a prompt consisting of incorrectly-answered questions in $m$, the chain-of-thought produced by the solver LLM, their incorrect predictions and the ground-truth answers. We ask the expert to provide one feedback for the mini batch (prompt is provided in Appendix A.12). The expert can suggest the following edits: add a section or subsection, delete a section or subsection, and edit a section or subsection.
Given the different edits for mini-batches within a batch $b$, we invoke the expert LLM again to summarize these edits into a single section update. This combination ensures some degree of smoothness at every update which helps stabilize training. To make sure that the expert is able to generate generalizable edits, we additionally provide a random set of incorrect examples that are not in the current batch and ask it to suggest an edit based on the existing edits that can correct the errors. As before, the class of edits allowed is the same.
History for effective exploration. To ensure comprehensive, non-repetitive exploration of prompts, we also provide the batch-level history of edits (Hu et al., 2024; Yang et al., 2023) in the mini-batch-level prompt. History $H[b]$ is presented as $\{e_{i},acc_{i}-acc_{i-1}\}$ where $e_{i}$ is the edit proposed at the $i^{th}$ update and $acc_{i}$ is the accuracy of the $i^{th}$ updated prompt (See Appendix A.12 for the prompt).